Question 3. A tire manufacturer warranties its tires to last at least 20,000 miles or “you get a new set of tires.” In its experience, a set of these tires lasts on average 26,000 miles with a standard deviation of 5,000 miles. Assume that the wear is normally distributed. The manufacturer profits $200 on each set sold, and replacing a set costs the manufacturer $400.

(a) What is the probability that a set of tires wears out before 20,000 miles?

(b) What is the probability that the manufacturer turns a profit on selling a set to one customer?

(c) If the manufacturer sells 500 sets of tires, what is the probability that it earns a profit after paying for any replacements? Assume that the purchases are made around the country and that the drivers experience independent amounts of wear.

The following content is partner provided
1
Expert's answer
2020-12-20T15:43:06-0500

μ = 26000

x = 20000

σ = 5000

(a)

"P(x<20000) = P(Z = \frac{x-μ}{σ}) \\

= P(Z = \frac{20000-26000}{5000}) \\

=P(Z = \frac{-6000}{5000}) \\

= P(Z=-1.2) \\

= 0.1151"

The probability that a set of tires wears out before 20,000 miles is 0.1151.

(b)

"P(x≥20000) = 1 – P(x<20000) \\

= 1 - P(Z = \frac{x-μ}{σ}) \\

= 1 - P(Z = \frac{20000-26000}{5000}) \\

= 1 - P(Z = \frac{-6000}{5000}) \\

= 1 - P(Z=-1.2) \\

= 1 - 0.1151 \\

= 0.8849"

The probability that the manufacturer turns a profit on selling a set to one customer is 0.8849

(c)

"P(x≥20000) = 1 – P(x<20000) \\

P(x<20000) = P(Z= \frac{x-μ}{\frac{σ}{\sqrt{n}}}) \\

P(x≥20000) = 1 - P(Z= \frac{x-μ}{\frac{σ}{\sqrt{n}}}) \\

n= 500 \\

P(x≥20000) = 1 - P(Z= \frac{20000-26000}{\frac{5000}{\sqrt{500}}}) \\

= 1 - P(Z = \frac{-6000}{226.61}) \\

= 1-P(Z=-26.832) \\

= 1-0 \\

=1"

The probability that it earns a profit after paying for any replacements is 1.